ES216-Homework on Dimensional Analysis 1. List the primary dimensions of (a). Velocity (b). Dynamics viscosity (c). Kinematic viscosity (d). Pressure
Answer #1

(a)

V=\frac{L}{t}

here L is the distance covered having primary dimension L and t is time taken to cover the distance having primary dimension T which has two different and independent primary dimensions. so, bvelocity is not a primary dimension

(b)Dynamic viscosity = M1L-1T-2 ( T1)

M1L-1T-1

(c)

v=\frac{\mu }{\rho }

v=\frac{\frac{m}{Lt}}{\frac{m}{L^{3}}}

v=\frac{L^{2}}{t}

(d)

(M/L·T2)

L2/t